# ConLID: Supervised Contrastive Learning for Low-Resource Language Identification

### v1

@@ 5.2 Fineweb-2 Analysis
[…]
@@
Resource ${|L|}$ ${|D|}$ Macro Micro
@@
In §5.1, we demonstrated that the incorporation of SCL was particularly beneficial for low-resource languages. To illustrate this effect, Table 5 presents the agreement scores for different language groups in Fineweb-2. Notably, low-resource languages exhibit a high degree of disagreement, with a micro agreement rate of $\sim$58.61%. While it is not feasible to manually verify individual cases of disagreement, it is crucial to emphasize that a 1% difference in low-resource language performance corresponds to $\sim$28,000 documents. This underscores the practical significance of even minor improvements in low-resource language identification (LID) models. As highlighted in §5.1, our findings confirm that the use of SCL effectively enhances performance for these languages.
@@ 6 Under-performing languages
We categorize languages based on domain availability and the number of sentences available in GlotLID-C-train-down. Figure 2 presents the corresponding F1 score performance on the UDHR dataset. We observe that models tend to exhibit improved generalization (i.e., higher F1 scores on UDHR) in languages for which they had a greater number of training domains. Models tend to generalize worse on languages for which training data was only available in one domain. Interestingly, while a majority of single-domain languages only had training data for the Bible domain, the model still achieves relatively high F1 scores for these languages, suggesting that training only on Bible data is not a solitary factor contributing to low performance on UDHR. Finally, we observe that the size of the training data does not appear to be a sole determinant of low performance on UDHR either.
Figure 3: UDHR misclassification pairs (languages with F1 $<$0.8) using ConLID-S. Most misclassification for a language happens between less than 5 languages.
@@
Our analysis reveals that most languages are predominantly misclassified as only a small subset of the 2099 total languages (i.e., number of unique misclassified languages < 10). This suggests that these languages share strong linguistic similarities with a limited set of other languages. Examples include Nyemba and Mbunda, both Bantu languages, and Northeastern and Southwestern Dinka, which exhibit significant linguistic overlap. Even languages with a high overall misclassification rate (number of unique misclassified languages > 10) tend to have one or a few dominant misclassification sources (i.e., highest misclassified language ratio), as seen in the case of Chimborazo Highland Quichua and Southern Pastaza Quechua.
@@ 7 Conclusion
In this work, we propose a novel approach that leverages supervised contrastive learning (SCL) for language identification (LID). The SCL objective explicitly encourages textual representations of the same language to form compact clusters while pushing apart representations of different languages in the embedding space. We evaluate our approach on three benchmark datasets: GlotLID, FLORES, and UDHR, demonstrating that incorporating SCL leads to substantial improvements over conventional cross-entropy (CE)-based LID methods. Specifically, our best SCL-based models achieve a 3.2% improvement over CE-based models for low-resource languages and a 6.8% improvement for languages with data from diverse domains, highlighting improved domain generalization, and underscoring the effectiveness of SCL in enhancing LID performance, particularly in challenging low-resource and cross-domain scenarios.
@@
Contrastive learning has been shown to enhance model performance and generalization by capturing a broader range of factors of variation compared to traditional classification networks. However, its effectiveness relies on access to large amounts of high-quality data, which poses a significant challenge in the development of Language Identification (LID) systems, where such data is often scarce.
Additionally, contrastive learning is highly sensitive to the selection of positive and negative examples, making the training process more complex. As a result, extensive ablation studies and careful hyperparameter tuning are necessary to optimize its performance and ensure its practical applicability.
While our training and in-domain test datasets encompass 2,099 languages, our out-of-domain evaluation dataset (UDHR) is limited to only 360 languages. This restriction constrains our ability to conduct a comprehensive analysis across the remaining languages.
Acknowledgements
We thank Deniz Bayazit, Ayush Kumar Tarun, María Grandury, Elif Kurtay, and Angelika Romanou for their helpful discussions and feedback on the draft. We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and a Meta LLM Evaluation Research Grant. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps.
@@ B.2 Hyperparameters
Appendix C Additional Results
C.1 UND labels recovery
@@
C.2 Script Analysis Break-down
We showed in §5.1 that while low-resource scripts have shown significant improvement over the baseline using ConLID-S, some high-resource scripts had more modest improvement. We show the break-down of the script analysis by the resource level in Table 16.
Script Resource Level #UDHR Languages $\text{LID}_{\text{CE}}$ $\Delta$$\text{LID}_{\text{SCL}}$ $\Delta$ConLID-S ConLID-S+LIDCE
Arabic Low 5 0.5176 0.1275 0.2824 0.1419
Arabic High 1 0.9388 0.0078 0.0249 0.0106
Cyrillic Low 5 0.9525 0.0072 0.0019 0.0038
Cyrillic High 24 0.9877 0.0039 0.0005 0.0023
Hani High 4 0.7570 -0.0468 -0.0088 0.0014
Java Low 1 0.8829 0.0923 0.1090 0.1007
Latin Low 20 0.8394 0.0039 0.0128 0.0108
Latin High 262 0.9560 -0.0037 0.0026 0.0025
Tifinagh Low 1 0.7158 0.0134 0.0134 0.0134
Tibetan High 2 0.9411 0.0253 0.0084 0.0084
Table 16: F1 score change of the LID models’ with SCL by language script broken down by the resource level. $\text{LID}_{\text{CE}}$ is used as a baseline, and subtracted from each model’s score to compute the absolute $\Delta$ score, e.g. $\Delta$$\text{LID}_{\text{SCL}}$= $\text{LID}_{\text{SCL}}$-$\text{LID}_{\text{CE}}$.
C.3 Analysis on Bible data
As noted GlotLID-C-train-down contains the following 5 domains: Bible, Literature, Politics, Multi-domain, Random. To test the effectiveness of our method for an extreme case, we performed the following experiment: We kept only the Bible, discarding the other domains from GlotLID-C-train-down, and trained 2 models with SCL and CE (ConLID-S-B, $\text{LID}_{\text{CE}}-\text{B}$). We, then, evaluated these 2 models on all 5 domains in addition to Flores-200 and UDHR. The macro F1-scores are reported in Table 17.
The results indicate that even with training is restricted to a single domain (i.e. Bible), our method enhances generalization to unseen domains. While our method yields modest improvements with a single domain exposure, the benefits are more pronounced when training on multiple domains, as shown in our analysis (§5).
Bible ID High 1400 0.9961 0.9961
Low 368 0.9883 0.9885
Random OOD High 202 0.5933 0.6075
Low 104 0.6429 0.6481
Multi-Domain OOD High 152 0.7996 0.8067
Low 90 0.7514 0.7444
Literature OOD High 119 0.8521 0.8638
Low 51 0.7641 0.7890
Government OOD High 5 0.9233 0.9137
Low 5 0.6263 0.6358
UDHR OOD High 246 0.8854 0.8886
Low 83 0.7961 0.8039
FLORES-200 OOD High 101 0.8938 0.8969
Low 66 0.8367 0.8375

### v2

@@ 5.3 FineWeb-2 Analysis
$\text{Agreement}=\frac{|\text{Pred. \textsubscript{\text{ConLID-S}}}\cap\text{Pred. \textsubscript{GlotLID-M{}}}|}{\text{\# of documents}}$
@@
Resource |L|{|L|} |D|{|D|} Macro Micro
@@
In §5.1, we demonstrated that the incorporation of SCL was particularly beneficial for low-resource languages. To illustrate its effect, Table 6 presents the agreement scores for different language groups in FineWeb-2. As expected, the agreement is high for high-resource languages, but much lower for low-resource languages, with a micro agreement rate of $\sim$58.61%. Since the dataset lacks ground-truth labels, it is not possible to directly determine which model is correct in cases of disagreement. However, as shown earlier in §5.1, incorporation of SCL outperforms the CE-based method, including GlotLID-M, on low-resource languages, suggesting that a low agreement score in these languages likely reflects ConLID-S making the correct predictions where the GlotLID-M fails. Importantly, even small differences are practically meaningful: a 1% improvement in low-resource language performance corresponds to $\sim$28,000 documents. This highlights the real-world impact of even minor improvements in low-resource LID models.
@@ 6 Under-performing languages
We categorize languages based on domain availability and the number of sentences available in GlotLID-C-train-down. Figure 2 presents the corresponding F1 score performance on the UDHR dataset. We observe that models tend to exhibit improved generalization (i.e., higher F1 scores on UDHR) in languages for which they had more diverse training domains. Models tend to generalize worse on languages for which training data was only available in one domain. Interestingly, while a majority of single-domain languages only had training data for the Bible domain, the model still achieves relatively high F1 scores for these languages, suggesting that training only on Bible data is not a sole factor contributing to low performance on UDHR. Finally, we observe that the size of the training data does not appear to be a sole determinant of low performance on UDHR either.
Figure 3: UDHR misclassification pairs (languages with F1 <<0.8) using ConLID-S. Most misclassification for a language happens between less than 5 languages.
@@
Our analysis reveals that most languages are predominantly misclassified as only a small subset of the 2099 total languages (i.e., number of unique misclassified languages < 10). This suggests that these languages share strong linguistic similarities with a limited set of other languages. Examples include Nyemba and Mbunda, both Bantu languages, and Northeastern and Southwestern Dinka, which exhibit significant linguistic overlap. Even languages with a high overall misclassification rate (number of unique misclassified languages > 10) tend to have one or a few dominant misclassification sources (i.e., highest misclassified language ratio), as seen in the case of Chimborazo Highland Quichua and Southern Pastaza Quechua. More examples of such pairs are provided in Table 17.
@@ 7 Conclusion
In this work, we propose a novel approach that leverages supervised contrastive learning (SCL) for language identification (LID). The SCL objective explicitly encourages textual representations of the same language to form compact clusters while pushing apart representations of different languages in the embedding space. We evaluate our approach on three benchmark datasets: GlotLID, FLORES, and UDHR, demonstrating that incorporating SCL leads to considerable improvements over conventional cross-entropy (CE)-based LID methods. Specifically, our best SCL-based models achieve a 3.2 percentage points improvement over CE-based models for low-resource languages and a 6.8 percentage points improvement for languages with data from diverse domains, highlighting improved domain generalization, and underscoring the effectiveness of SCL in enhancing LID performance, particularly in challenging low-resource and cross-domain scenarios. Moreover, we also demonstrate that SCL and CE-based methods are complementary to each other and can be used together during inference, as in the case of GlotLID-M+ConLID-S.
@@
Contrastive learning has been shown to enhance model performance and generalization by capturing a broader range of factors of variation compared to traditional classification networks. However, its effectiveness relies on access to high-quality data from diverse domains, which poses a challenge in the development of Language Identification (LID) systems, where such data is often scarce.
While our training and in-domain test datasets encompass 2,099 languages, our out-of-domain evaluation dataset (UDHR) is limited to only 360 languages. This restriction constrains our ability to conduct a comprehensive analysis across the remaining languages. Additionally, the lack of evaluation dataset in other domains limits our ability to thoroughly assess method’s generalization capacity.
Acknowledgments
We thank Deniz Bayazit, Angelika Romanou, and Ayush Kumar Tarun for their helpful discussions and feedback on our paper. We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), the European Research Council (Starting grant no. 101222478, RESPECT-LM), the AI2050 program at Schmidt Sciences (Grant #G-25-69783), Sony Group Corporation, and the Swiss National Supercomputing Center (CSCS) in the form of an infrastructure engineering and development project.
@@ B.2 Hyperparameters
Appendix C Hard Selection Algorithm
The negatives pairs in the Hard selection method, as explained in §3.3, are sampled according to the Algorithm 1.
Algorithm 1 Hard Negative Sample Selection
1:for $i<B$ do
3:  if $|N(i)|<K$ then
4:   $N(i)=\left\{z_{j}\in S,y_{j}\neq y_{i},s_{j}=s_{i}\right\}$ {Condition 2}
5:  end if
6:  if $|N(i)|<K$ then
7:   $N(i)=\left\{z_{j}\in S,y_{j}\neq y_{i},d_{j}=d_{i}\right\}$ {Condition 3}
8:  end if
9:  if $|N(i)|<K$ then
10:   $N(i)=\left\{z_{j}\in S,y_{j}\neq y_{i}\right\}$ {Condition 4}
11:  end if
12:end for
13:return $N(i)$ {negative pairs for each sample}
$L_{1}$ $L_{2}$ # errors
krl_Latn olo_Latn 183 (100%)
bam_Latn dyu_Latn 8 (100%)
hrv_Latn hbs_Latn 45 (98%)
srp_Latn hbs_Latn 29 (88%)
hak_Hani wuu_Hani 43 (77%)
wuu_Hani cmn_Hani 3 (75%)
qvn_Latn quy_Latn 23 (64%)
yue_Hani cmn_Hani 8 (62%)
zgh_Tfng tzm_Tfng 14 (54%)
fuf_Latn fuv_Latn 8 (27%)
Table 17: Examples of language pairs where the first language is frequently misclassified as the second.
Appendix D Additional Results
Bible ID High 1400 0.9961 0.9961
Low 368 0.9883 0.9885
Random OOD High 202 0.5933 0.6075
Low 104 0.6429 0.6481
Multi-Domain OOD High 152 0.7996 0.8067
Low 90 0.7514 0.7444
Literature OOD High 119 0.8521 0.8638
Low 51 0.7641 0.7890
Government OOD High 5 0.9233 0.9137
Low 5 0.6263 0.6358
UDHR OOD High 246 0.8854 0.8886
Low 83 0.7961 0.8039
FLORES-200 OOD High 101 0.8938 0.8969
Low 66 0.8367 0.8375
D.1 Significance testing using bootstrapping
We additionally applied a t-test using bootstrapping as a significance testing method to more clearly demonstrate the effectiveness of SCL approach (Table 19). Based on the bootstrapped results, p-value is less than 0.05, and the SCL method still outperforms both the GlotLID-M and the cross-entropy baseline, $\text{LID}_{\text{CE}}$.
The relatively lower scores for the UDHR dataset can be attributed to the imbalance in the number of sentences across languages. Some overrepresented languages exhibit lower performance, and their frequent selection during bootstrapping iterations negatively impacts the overall results.
GlotLID-C Test UDHR FLORES-200
$\text{LID}_{\text{CE}}$ 0.9874 0.6080 0.8598
ConLID-S 0.9878 0.6176 0.8616
GlotLID-M 0.9877 0.5771 0.8614
GlotLID-M+ConLID-S 0.9943 0.6201 0.8859
Table 19: Macro F1 score via bootstrapping with 1000 iterations (p<0.05)
D.2 Script Analysis Breakdown
We showed in §5.1 that while low-resource scripts have shown significant improvement over the baseline using ConLID-S, some high-resource scripts had more modest improvement. We show the break-down of the script analysis by the resource level in Table 20. We attribute the performance gains primarily to the resource level of the languages associated with each script.
While the number of low-resource languages is typically smaller, those within a given script tend to show greater improvements in performance compared to their high-resource counterparts. Additionally, the number of languages per script varies significantly. For instance, the Latin script encompasses 282 languages in this table, whereas the substantial improvement seen in the Java script is based on a single language. This discrepancy in language count per script also contributes to the observed performance variance.
[…]
Arabic Low 5 0.5176 0.1275 0.2824 0.1419
Arabic High 1 0.9388 0.0078 0.0249 0.0106
Cyrillic Low 5 0.9525 0.0072 0.0019 0.0038
Cyrillic High 24 0.9877 0.0039 0.0005 0.0023
Hani High 4 0.7570 -0.0468 -0.0088 0.0014
Java Low 1 0.8829 0.0923 0.1090 0.1007
Latin Low 20 0.8394 0.0039 0.0128 0.0108
Latin High 262 0.9560 -0.0037 0.0026 0.0025
Tifinagh Low 1 0.7158 0.0134 0.0134 0.0134
Tibetan High 2 0.9411 0.0253 0.0084 0.0084
[…]
D.3 Resource level categorization
Languages can be categorized as low- or high-resource based on either the availability of large amounts of unlabeled data (e.g., pretraining data collected from the web) or the quantity of labeled data available for specific tasks. In many cases, these two criteria coincide. However, in this study, our primary goal is to evaluate the effectiveness of our approach under conditions where labeled data for training an LID system is scarce. For this reason, we adopt the second criterion — the amount of labeled training data — as the basis for our categorization. Specifically, we classify languages with fewer than 10,000 labeled examples as low-resource, and those with 10,000 or more labeled examples as high-resource, as described in Section 5.1. We find this definition particularly meaningful because the performance of language models is strongly influenced by the amount of labeled data available for training. To provide a complementary perspective, we also analyze our results using the categorization proposed by Joshi et al. (2020b). Table 21 presents the performance of our method on the UDHR dataset according to this scheme. Labels 0 through 5 follow the definitions from Joshi et al. (2020b), and we introduce an additional label, -1, for languages not included in their original categorization (low-resource and under-represented languages). As shown, most of the performance gains occur within these labeled groups, highlighting the robustness of our approach across different categorization schemes.
Language Class $|L|$ $\text{LID}_{\text{CE}}$ ConLID-S
-1 135 0.8183 0.0204
0 93 0.9059 0.0004
1 76 0.9554 0.0036
2 13 0.9686 -0.0024
3 20 0.9830 0.0067
4 17 0.9311 -0.0020
5 6 0.9812 0.0008
Table 21: The performance of the models on UDHR dataset based on the categorization low- and high-resource languages of Joshi et al. (2020b).
D.4 UND labels recovery